# Null projections and noncommutative function theory in operator algebras

### v1

@@ Abstract.
We study projections in the bidual of a $\mathrm{C}^{*}$-algebra $B$ that are null with respect to a subalgebra $A$, that is projections $p\in B^{**}$ satisfying $|\varphi|(p)=0$ for every $\varphi\in B^{*}$ annihilating $A$. In the separable case, $A$-null projections are precisely the peak projections in the bidual of $A$ at which the subalgebra $A$ interpolates the entire $\mathrm{C}^{*}$-algebra $B$. These are analogues of null sets in classical function theory, on which several profound results rely. This motivates the development of a noncommutative variant, which we use to find appropriate ‘quantized’ versions of some of these classical facts. Through a delicate generalization of a theorem of Varopoulos, we show that, roughly speaking, sufficiently regular interpolation projections are null precisely when their atomic parts are. As an application, we give alternative proofs and sharpenings of some recent peak-interpolation results of Davidson and Hartz for algebras on Hilbert function spaces, also illuminating thereby how earlier noncommutative peak-interpolation theory may be applied. In another direction, given a convex subset of the state space of $B$, we characterize when the associated Riesz projection is null. This is then applied to various important topics in noncommutative function theory, such as the F.& M. Riesz property, the existence of Lebesgue decompositions, the description of Henkin functionals, and Arveson’s noncommutative Hardy spaces (maximal subdiagonal algebras).
@@ 1. Introduction
Let $B$ be a unital $\mathrm{C}^{*}$-algebra. A projection $u$ in the von Neumann algebra $B^{**}$ is open if it is the increasing weak-*** limit of a net of positive contractions in $B$. Likewise, a projection $q\in B^{**}$ is closed if $1-q$ is open. These notions of non-commutative topology were introduced and developed by Akemann [akemann1969, akemann1970left].
@@
When $q$ satisfies that $Aq=Bq$ as subspaces in $B^{**}$, we say that $q$ is an $A$-interpolation projection. Note that this is a one-sided concept, in that it is not equivalent to the equality $qA=qB$ in general (as one may see even in the upper triangular $2\times 2$ matrices). When $B$ is commutative, closely related notions of interpolation have long been of interest to function theorists, even for countable sets (see for instance [AHMR2019interp] and the references therein).
@@
(i) $\Rightarrow$(ii): (This does not require that $A$ be a subalgebra.) That $q\in A^{\perp\perp}$ follows from the inequality $|\varphi(q)|^{2}\leq|\varphi|(q)\,|\varphi|(1)$ if $\varphi\in A^{\perp}$ (see Lemma 2.1 below). That $Aq=Bq$ is proved in the unital case in the first lines of the proof of [hay2007, Proposition 3.4]. The nonunital case follows by the same proof (which proof is given with more detail in the proof that (ii) implies (iii) in [CT2023Henkin, Theorem 6.2]) once we know that $Aq$ is closed.
To see this, we simply need to adapt the proof of [hay2007, Proposition 3.1] to the case where $X=A$ and $B$ is nonunital. Consider $B^{1}=B+{\mathbb{C}}1\subset B^{**}$. Following the argument in [hay2007], we find a decreasing net $(e_{t})$ in $B^{1}$ converging to $q$ in the weak-*** topology of $B^{**}$. While $e_{t}$ may not belong to $B$, we nevertheless have $e_{t}B\subset B$ for each index $t$, which is all that is needed in Hay’s argument to conclude that $Aq$ is isometric to the quotient of $A$ by some norm-closed subspace, and hence is closed. ∎
@@
Section 4 is concerned with our noncommutative version of Varopoulos’ theorem. For this purpose, we need the following “complete” refinement of the notion of interpolation projection. For $\kappa>0$, we say that a projection $q$ is a complete $(A,\kappa)$-interpolation projection if, given any $n\geq 1$ and any matrix $[b_{ij}]\in{\mathbb{M}}_{n}(B)$, there are matrices $[a_{ij}],[a^{\prime}_{ij}]\in{\mathbb{M}}_{n}(A)$ such that
@@
We interpret Theorem A as saying that for sufficiently ‘topologically regular’ interpolation projections (i.e. where such $u$ may approach $q$), the property of being $A$-null is entirely captured by the atomic part. In Proposition 4.6, we exhibit a non-trivial class of algebras where the atomic part of every projection lies in $A^{\perp\perp}$. In Corollary 4.7, we identify some natural sufficient conditions under which the conclusion of Theorem A can be strengthened to $q$ being null. In particular, we show in Corollary 4.9 how to recover Varopoulos’ original theorem using our results.
@@
We omit the precise definitions for now, and refer the reader to the relevant section for more detail. Let ${\mathcal{H}}$ be a Hilbert space of analytic functions on the open unit ball ${\mathbb{B}}_{d}\subset{\mathbb{C}}^{d}$, admitting a reproducing kernel that is unitarily invariant and regular. There is a central projection $z\in{\mathcal{T}}({\mathcal{H}})^{**}$ and a natural ***-isomorphism $\Theta:{\mathcal{T}}({\mathcal{H}})^{**}z\to\mathrm{C}({\mathbb{S}}_{d})^{**}$.
@@
Beyond this result, our approach can be interpreted as general method that we hope will be applicable in other contexts to prove strong peak-interpolation results.
In Section 6, we turn to a different problem regarding $A$-null projections. Let $\Delta$ be a norm-closed convex set of states on $B$. We let $\operatorname{AC}(\Delta)\subset B^{*}$ denote the set of those functionals that are absolutely continuous with respect to some element of $\Delta$. It is known [CT2023Henkin, Theorem 3.5] that there is a unique projection $r_{\Delta}\in B^{**}$ with the property that $**\operatorname{AC}(\Delta)^{\perp}=(1-r_{\Delta})B^{**}$. We call this projection the Riesz projection corresponding to $\Delta$. We aim to characterize when this Riesz projection satisfies that $1-r_{\Delta}$ is $A$-null.
More precisely, our main result relates this condition to other natural properties inspired by classical results in function theory. We say that $A$ has the F.&M. Riesz property in $B$ with respect to $\Delta$ if $\varphi(r_{\Delta}\cdot)\in A^{\perp}$ for every $\varphi\in A^{\perp}$. This property has been a recurring theme (under various guises) in the operator algebras literature; see for instance [BL2007, CH2022] and the references therein.
In addition, we say that $A$ has the Forelli property in $B$ with respect to $\Delta$ if for every projection $q\in\operatorname{AC}(\Delta)^{\perp}$ that is the supremum of a collection of closed projections, there is a contractive net $(a_{i})$ in $A$ such that $(qa_{i})$ converges to 00 in the weak-*** topology of $B^{**}$, and $(\varphi(1-a_{i}))$ converges to 00 for every $\varphi\in\operatorname{AC}(\Delta)$. This technical approximation property is directly inspired by [rudin2008, Lemma 9.5.5]. Our next main result can be streamlined as follows.
@@
Assume that $\Delta$ is closed in the weak-*** topology of $B^{**}$. Consider the following statements.
@@
    $A$ has the Forelli property in $B$ with respect to $\Delta$.
@@
    $A$ has the F.&M. Riesz property in $B$ with respect to $\Delta$.
@@
The condition that $\Delta$ be weak-*** closed is used in proving Lemma 6.4, a non-commutative analogue of a classical lemma due to Rainwater [rainwater1969]. Without it, the topological nature of the Riesz projection is not as clear. Nevertheless, many of the implications above remain valid if $\Delta$ is merely assumed to be norm-closed (see Theorem 6.5). We apply these findings in Section 7 to the existence of Lebesgue decompositions for operator algebras, which is a main theme in the recent paper of the second author and Hartz [CH2022]. In Corollary 7.4, we give a new sufficient condition for such decompositions to exist.
@@
In the last two sections of the paper, we aim to understand when all the statements in Theorem B are in fact equivalent. Our first approach, used in Section 8, consists of comparing the F.& M. Riesz property defined above with a stronger variant of it. The key result in this direction is Corollary 8.2, showing that the so-called strong F.& M. Riesz property can be reformulated in terms of a band property. This is in fact a consequence of a certain rigidity phenomenon (Theorem 8.1) resulting from the F.& M. Riesz property, and which is related to the classical Cole-Range theorem [rudin2008, Theorem 9.6.1]. This in turn is shown in Examples 6 and 7 to unify some previously known results on the structure of Henkin functionals [CD2016duality, CH2022, CT2023Henkin]. The property of the set of Henkin functionals forming a band plays an important role in these sections. For example our analysis shows that much of the function theory developed in Chapter 9 of [rudin2008], and in particular the Cole-Range theorem, rests on two main pillars: that the Henkin functionals form band, and the F.& M. Riesz property.
@@
Finally, in Section 9, we tackle the equivalence question for Theorem B using a different route, centered around the following notion. We say that $A$ has the Amar–Lederer property in $B$ with respect to $\Delta$ if all closed projections in $\operatorname{AC}(\Delta)^{\perp}$ are dominated by a closed projection in $\operatorname{AC}(\Delta)^{\perp}\cap A^{\perp\perp}$. A very important classical result of Amar–Lederer [AL1971] shows that $H^{\infty}({\mathbb{D}})$ satisfies this property inside $L^{\infty}({\mathbb{T}})$ with respect to the state of integration against arclength measure: speaking slightly loosely they show that any Lebesgue null Borel set is contained in a (closed) Lebesgue null peak set. On the other hand, the statements in Theorem B fail to be equivalent in this case (Example 8). The rest of the section is devoted to proving Theorem 9.4, where we contribute to the theory of maximal subdiagonal algebras. Indeed, we refine known results from [ueda2009, BL2018] using some new noncommutative measure theory, and show that such algebras do satisfy the full Amar and Lederer property for all null closed (indeed ‘regular’) projections, and hence they also have the Forelli property.
@@ 2. Null projections
Let $B$ be a $\mathrm{C}^{*}$-algebra. Then, the compact projections in $B^{**}$ span a weak-*** dense subspace.
@@
Let $D\subset B^{**}$ denote the subspace generated by all compact projections in $B^{**}$. Let $\varphi\in B^{*}$ be a functional annihilating $D$. We claim that $\varphi=0$. To see this, it suffices to fix a positive element $b\in B$ and to show that $\varphi(b)=0$. Let $\Omega\subset{\mathbb{R}}$ denote the spectrum of $b$. We recall that $\mathrm{C}_{0}(\Omega\setminus\{0\})$ is ***-isomorphic to the $\mathrm{C}^{*}$-algebra generated by $b$. Then there exists a regular Borel measure $\mu$ on $\Omega\setminus\{0\}$ such that
[…]
for every $f\in\mathrm{C}_{0}(\Omega)$ with $f(0)=0$. If $K\subset\Omega\setminus\{0\}$ is compact then $\chi_{K}(b)\in B^{**}$ by functional calculus, and $\chi_{K}(b)$ is closed with respect to the unitization of $B$, hence it is a compact projection in $B^{**}$[AAP1989, Lemma 2.4]. Hence, $\chi_{K}(b)\in D$ and $\varphi(\chi_{K}(b))=0$. On the other hand, by a dominated convergence argument applied to a sequence of continuous functions decreasing to $\chi_{K}$ and vanishing at 00, we find by virtue of (1) that $\varphi(\chi_{K}(b))=\mu(K)$ so that $\mu(K)=0$. By regularity of the measure $\mu$, we conclude that $\mu=0$, and thus $\varphi=0$. ∎
@@
(i)$\Rightarrow$(ii): By assumption, we see that $M^{\perp}=\{0\}$, so that, vacuously, the unit of $B^{**}$ is $M$-null.
@@
(iii)$\Rightarrow$(i): The assumption combined with Lemma 2.2 implies that $M^{\perp\perp}=B^{**}$. In turn, by the Hahn–Banach theorem, we find $M=B$ since $M$ is norm-closed. ∎
@@
Let $B$ be a unital $\mathrm{C}^{*}$-algebra and let $q\in B^{**}$ be a closed projection. Denote by $B_{q}\subset B$ the $\mathrm{C}^{*}$-subalgebra of elements in $B$ commuting with $q$. We say that a projection $r\in B_{q}^{\perp\perp}$ is closed with respect to $B_{q}$ if there is a net $(b_{i})$ of positive contractions in $B_{q}$ decreasing to $r$ in the weak-*** topology of $B^{**}$. Equivalently, using the weak-*** homeomorphic ***-isomorphism $B_{q}^{**}\cong B_{q}^{\perp\perp}$, this means that $r$ is closed when viewed as an element of $B_{q}^{**}$. It is easily verified that $q$ itself lies in $B_{q}^{\perp\perp}$, and it is closed with respect to $B_{q}$ (see for instance [blecher2022, Lemma 2.1]). Indeed a projection $r\in B_{q}^{\perp\perp}$ is closed with respect to $B_{q}$ if and only if it is closed with respect to $B$.
@@
    $q$ is $A_{q}$-null with respect to $B_{q}$.
@@
    $q$ is $A_{q}$-interpolating with respect to $B_{q}$.
@@
By Lemma 2.1 in [blecher2022] we automatically have $q\in A_{q}^{\perp\perp}$. Then the equivalence of (ii) and (iii) follows from the “Basic principle” stated in the introduction.
@@
(ii)$\Rightarrow$(i): Let $p\in B^{**}$ be any projection lying in $B_{q}^{\perp\perp}$ and dominated by $q$. Then, $p$ may be viewed as a projection in $B_{q}^{**}$ dominated by $q$. Our assumption then implies that $|\varphi|(p)=0$ for every $\varphi\in B_{q}^{*}$ annihilating $A_{q}$. In particular, by virtue of Lemma 2.1, we infer that $\varphi(p)=0$ for all such functionals $\varphi$, and hence $p$ lies in the double annihilator of $A_{q}$ inside of $B_{q}^{**}$. In turn, when viewed as a projection in $B^{**}$, this means that $p$ lies in $A_{q}^{\perp\perp}$. ∎

### v2

@@ Abstract.
We study projections in the bidual of a $\mathrm{C}^{*}$-algebra $B$ that are null with respect to a subalgebra $A$, that is projections $p\in B^{**}$ satisfying $|\varphi|(p)=0$ for every $\varphi\in B^{*}$ annihilating $A$. In the separable case, $A$-null projections are precisely the peak projections in the bidual of $A$ at which the subalgebra $A$ interpolates the entire $\mathrm{C}^{*}$-algebra $B$. These are analogues of null sets in classical function theory, on which several profound results rely. This motivates the development of a noncommutative variant, which we use to find appropriate ‘quantized’ versions of some of these classical facts. Through a delicate generalization of a theorem of Varopoulos, we show that, roughly speaking, sufficiently regular interpolation projections are null precisely when their atomic parts are. As an application, we give alternative proofs and sharpenings of some recent peak interpolation results of Davidson and Hartz for algebras on Hilbert function spaces, also illuminating thereby how earlier noncommutative peak-interpolation theory may be applied. In another direction, given a convex subset of the state space of $B$, we characterize when the associated Riesz projection is null. This is then applied to various important topics in noncommutative function theory, such as the F.& M. Riesz property, the existence of Lebesgue decompositions, the description of Henkin functionals, and Arveson’s noncommutative Hardy spaces (maximal subdiagonal algebras).
2000 Mathematics Subject Classification:
46L52, 46L51, 47L30
@@ 1. Introduction
Let $B$ be a unital $\mathrm{C}^{*}$-algebra. A projection $u$ in the von Neumann algebra $B^{**}$ is open if it is the increasing weak-* limit of a net of positive contractions in $B$. Likewise, a projection $q\in B^{**}$ is closed if 1q1-q is open. These notions of non-commutative topology were introduced and developed by Akemann [akemann1969, akemann1970left].
@@
When $q$ satisfies that $qA=qB$ as subspaces in $B^{**}$, we say that $q$ is an $A$-interpolation projection. Note that this is a one-sided concept, in that it is not equivalent to the equality $Aq=Bq$ in general (as one may see even in the upper triangular $2\times 2$ matrices). When $B$ is commutative, closely related notions of interpolation have long been of interest to function theorists, even for countable sets (see for instance [AHMR2019interp] and the references therein).
@@
(i) $\Rightarrow$(ii): (This does not require that $A$ be a subalgebra.) That $q\in A^{\perp\perp}$ follows from the inequality $|\varphi(q)|^{2}\leq|\varphi|(q)\,|\varphi|(1)$ if $\varphi\in A^{\perp}$ (see Lemma 2.1 below). That $qA=qB$ is proved in the unital case in the first lines of the proof of [hay2007, Proposition 3.4]. The nonunital case follows by the same proof (which proof is given with more detail in the proof that (ii) implies (iii) in [CT2023Henkin, Theorem 6.2]) once we know that $qA$ is closed.
To see this, we simply need to adapt the proof of [hay2007, Proposition 3.1] to the case where $X=A$ and $B$ is nonunital. Consider $B^{1}=B+{\mathbb{C}}1\subset B^{**}$. Following the argument in [hay2007], we find a decreasing net $(e_{t})$ in $B^{1}$ converging to $q$ in the weak-* topology of $B^{**}$. While $e_{t}$ may not belong to $B$, we nevertheless have $Be_{t}\subset B$ for each index $t$, which is all that is needed in Hay’s argument to conclude that $qA$ is isometric to the quotient of $A$ by some norm-closed subspace, and hence is closed. ∎
@@
Section 4 is concerned with our noncommutative version of Varopoulos’ theorem. For this purpose, we need the following “complete” refinement of the notion of interpolation projection. For $\kappa>0$, we say that a projection $q$ is a (two-sided) complete $(A,\kappa)$-interpolation projection if, given any $n\geq 1$ and any matrix $b_{ij}]\in{\mathbb{M}}_{n}(B)$, there are matrices $a_{ij}],[a^{\prime}_{ij}]\in{\mathbb{M}}_{n}(A)$ such that
@@
We interpret Theorem A as saying that for sufficiently ‘topologically regular’ interpolation projections (i.e. where such $u$ may approach $q$) commuting with $A$, the property of being $A$-null is entirely captured by the atomic part. In Proposition 4.6, we exhibit a non-trivial class of algebras where the atomic part of every projection lies in $A^{\perp\perp}$. In Corollary 4.7, we identify some natural sufficient conditions under which the conclusion of Theorem A can be strengthened to $q$ being null. In particular, we show in Corollary 4.9 how to recover Varopoulos’ original theorem using our results.
@@
We omit the precise definitions for now, and refer the reader to the relevant section for more detail. Let $\mathcal{H}$} be a Hilbert space of analytic functions on the open unit ball ${\mathbb{B}}_{d}\subset{\mathbb{C}}^{d}$, admitting a reproducing kernel that is unitarily invariant and regular. There is a central projection $z\in{\mathcal{T}}({\mathcal{H}})^{**}$ and a natural *-isomorphism $\Theta:{\mathcal{T}}({\mathcal{H}})^{**}z\to\mathrm{C}({\mathbb{S}}_{d})^{**}$.
@@
Beyond this result, our approach can be interpreted as a general method that we hope will be applicable in other contexts to prove strong peak-interpolation results.
In Section 6, we turn to a different problem regarding $A$-null projections. Let $\Delta$ be a norm-closed convex set of states on $B$. We let $\operatorname{AC}(\Delta)\subset B^{*}$ denote the set of those functionals that are absolutely continuous (see Section 6 for the definition of this) with respect to some element of $\Delta$. It is known [CT2023Henkin, Theorem 3.5] that there is a unique projection $r_{\Delta}\in B^{**}$ with the property that $\operatorname{AC}(\Delta)^{\perp}=(1-r_{\Delta})B^{**}$. We call this projection the Riesz projection corresponding to $\Delta$. We aim to characterize when this Riesz projection satisfies that $1-r_{\Delta}$ is $A$-null.
More precisely, our main result relates this condition to other natural properties inspired by classical results in function theory. We say that $A$ has the F.&M. Riesz property in $B$ with respect to $\Delta$ if $\varphi(r_{\Delta}\cdot)\in A^{\perp}$ for every $\varphi\in A^{\perp}$. This property has been a recurring theme (under various guises) in the operator algebras literature; see for instance [BL2007, CH2025] and the references therein.
In addition, we say that $A$ has the Forelli property in $B$ with respect to $\Delta$ if for every projection $q\in\operatorname{AC}(\Delta)^{\perp}$ that is the supremum of a collection of closed projections, there is a contractive net $(a_{i})$ in $A$ such that $(qa_{i})$ converges to 0 in the weak-* topology of $B^{**}$, and $(\varphi(1-a_{i}))$ converges to 0 for every $\varphi\in\operatorname{AC}(\Delta)$. This technical approximation property is directly inspired by [rudin2008, Lemma 9.5.5]. Our next main result can be streamlined as follows.
@@
Assume that $\Delta$ is closed in the weak-* topology of $B^{**}$. Consider the following statements.
@@
    AA has the Forelli property in $B$ with respect to $\Delta$.
@@
    AA has the F.&M. Riesz property in $B$ with respect to $\Delta$.
@@
The condition that $\Delta$ be weak-* closed is used in proving Lemma 6.4, a non-commutative analogue of a classical lemma due to Rainwater [rainwater1969]. Without it, the topological nature of the Riesz projection is not as clear. Nevertheless, many of the implications above remain valid if $\Delta$ is merely assumed to be norm-closed (see Theorem 6.5). We apply these findings in Section 7 to the existence of Lebesgue decompositions for operator algebras, which is a main theme in the recent paper of the second author and Hartz [CH2025]. In Corollary 7.4, we give a new sufficient condition for such decompositions to exist.
@@
In the last two sections of the paper, we aim to understand when all the statements in Theorem C are in fact equivalent. Our first approach, used in Section 8, consists of comparing the F.& M. Riesz property defined above with a stronger variant of it. The key result in this direction is Corollary 8.2, showing that the so-called strong F.& M. Riesz property can be reformulated in terms of a band property. This is in fact a consequence of a certain rigidity phenomenon (Theorem 8.1) resulting from the F.& M. Riesz property, and which is related to the classical Cole-Range theorem [rudin2008, Theorem 9.6.1]. This in turn is shown in Examples 6 and 7 to unify some previously known results on the structure of Henkin functionals [CD2016duality, CH2025, CT2023Henkin]. The property of the set of Henkin functionals forming a band plays an important role in these sections. For example our analysis shows that much of the function theory developed in Chapter 9 of [rudin2008], and in particular the Cole-Range theorem, rests on two main pillars: that the Henkin functionals form a band, and the F.& M. Riesz property. Also, our work on the (strong) F.& M. Riesz property may be viewed as generalizations of Bishop’s insight [bishop1962] that a (strong) F.& M. Riesz theorem implies a generalized Rudin-Carleson theorem (the latter corresponds for us to the appropriate projection(s) being $A$-null).
@@
Finally, in Section 9, we tackle the equivalence question for Theorem C using a different route, centered around the following notion. We say that $A$ has the Amar–Lederer property in $B$ with respect to $\Delta$ if all closed projections in $\operatorname{AC}(\Delta)^{\perp}$ are dominated by a closed projection in $\operatorname{AC}(\Delta)^{\perp}\cap A^{\perp\perp}$. A very important classical result of Amar–Lederer [AL1971] shows that $H^{\infty}({\mathbb{D}})$ satisfies this property inside $L^{\infty}({\mathbb{T}})$ with respect to the state of integration against arclength measure: speaking slightly loosely they show that any Lebesgue null Borel set is contained in a (closed) Lebesgue null peak set. On the other hand, the statements in Theorem C fail to be equivalent in this case (Example 8). The rest of the section is devoted to proving Theorem 9.4, where we contribute to the theory of maximal subdiagonal algebras. Indeed, we refine known results from [ueda2009, BL2018] using some new noncommutative measure theory, and show that such algebras do satisfy the full Amar and Lederer property for all null closed (indeed ‘regular’) projections, and hence they also have the Forelli property. We say ‘full’ because of a severe restriction imposed in the latter papers on the null projections used.
@@ 2. Null projections
Let $B$ be a $\mathrm{C}^{*}$-algebra. Then, the compact projections in $B^{**}$ span a weak-* dense subspace.
@@
Let $D\subset B^{**}$ denote the subspace generated by all compact projections in $B^{**}$. Let $\varphi\in B^{*}$ be a functional annihilating $D$. We claim that $\varphi=0$. To see this, it suffices to fix a positive element $b\in B$ and to show that $\varphi(b)=0$. Let $\Omega\subset{\mathbb{R}}$ denote the spectrum of $b$. We recall that $\mathrm{C}_{0}(\Omega\setminus\{0\})$ is *-isomorphic to the $\mathrm{C}^{*}$-algebra generated by $b$. Then there exists a regular Borel measure $\mu$ on $\Omega\setminus\{0\}$ such that
(1) $\varphi(f(b))=\int_{\Omega}fd\mu$
for every $f\in\mathrm{C}_{0}(\Omega)$ with $f(0)=0$. If $K\subset\Omega\setminus\{0\}$ is compact then $\chi_{K}(b)\in B^{**}$ by functional calculus, and $\chi_{K}(b)$ is closed with respect to the unitization of $B$, hence it is a compact projection in $B^{**}$[AAP1989, Lemma 2.4]. Hence, $\chi_{K}(b)\in D$ and $\varphi(\chi_{K}(b))=0$. On the other hand, by a dominated convergence argument applied to a sequence of continuous functions decreasing to $\chi_{K}$ and vanishing at 0, we find by virtue of (1) that $\varphi(\chi_{K}(b))=\mu(K)$ so that $\mu(K)=0$. By regularity of the measure $\mu$, we conclude that $\mu=0$, and thus $\varphi=0$. ∎
@@
[…]
@@
[…]
@@
Let $B$ be a unital $\mathrm{C}^{*}$-algebra and let $q\in B^{**}$ be a closed projection. Denote by $B_{q}\subset B$ the $\mathrm{C}^{*}$-subalgebra of elements in $B$ commuting with $q$. We say that a projection $r\in B_{q}^{\perp\perp}$ is closed with respect to $B_{q}$ if there is a net $(b_{i})$ of positive contractions in $B_{q}$ decreasing to $r$ in the weak-* topology of $B^{**}$. Equivalently, using the weak-* homeomorphic *-isomorphism $B_{q}^{**}\cong B_{q}^{\perp\perp}$, this means that $r$ is closed when viewed as an element of $B_{q}^{**}$. It is easily verified that $q$ itself lies in $B_{q}^{\perp\perp}$, and it is closed with respect to $B_{q}$ (see for instance [blecher2022, Lemma 2.1]). Indeed a projection $r\in B_{q}^{\perp\perp}$ is closed with respect to $B_{q}$ if and only if it is closed with respect to $B$.
@@
    $q$ is $A_{q}$-null with respect to $B_{q}$ (that is, $|\varphi|(q)=0$ for all $\varphi\in(A_{q})^{\perp}$).
@@
    $q$ is $A_{q}$-interpolating with respect to $B_{q}$ (that is, $qA_{q}=qB_{q}$).
@@
By Lemma 2.1 in [blecher2022] we automatically have $q\in A_{q}^{\perp\perp}$. Then the equivalence of (ii) and (iii) follows from the “Basic principle” stated in the introduction. Note that by definition of $B_{q}$ the projection $q$ commutes with the elements here.
@@
[…]